# Highly efficient quantum Stirling engine using multilayer Graphene

### v1

@@ III.2 Energy Scales and Parametrization of the Chemical Potential to Conserve Particle Number
Within the field range $0.1–5\,\mathrm{T}$, the energy separation between consecutive Landau levels in the trilayer remains consistently smaller than in the bilayer and well below that of the monolayer. This hierarchy implies that monolayer graphene requires substantially higher temperatures to thermally populate its first excited Landau level, whereas the trilayer can achieve full population at comparatively lower temperatures.
Since the thermodynamic quantities explicitly depend on the energy levels through the chemical potential, it is essential to examine how $\mu$ evolves with the magnetic field $B$. To this end, we study the total particle number $\langle N\rangle$, whose dependence on $B$ and $\mu$ provides indirect yet crucial information about the chemical potential. This behavior is shown in Fig. 2, which presents contour plots of $\langle N\rangle$ as a function of the magnetic field for monolayer (a), bilayer (b), and trilayer (c) graphene at a fixed temperature $T=50\,\mathrm{K}$.
In Fig. 2(a), the monolayer exhibits contours of $\langle N\rangle$ that approximately follow a square-root scaling, $\mu\propto\sqrt{B}$, consistent with the relativistic Landau-level spectrum $\varepsilon_{n}^{(1)}\propto\sqrt{nB}$ characteristic of Dirac fermions.
For the AB-stacked bilayer in Fig. 2(b), the contour lines reveal an almost linear dependence of $\mu$ on $B$, indicating a scaling of the form $\mu\propto B$ while preserving the particle number. This trend reflects the bilayer Landau-level structure $\varepsilon_{n}^{(2)}\propto B\sqrt{n(n-1)}$, where carriers behave as massive quasiparticles with a linear field dependence at low energies.
In the trilayer case, shown in Fig. 2(c), the scaling becomes significantly steeper. The chemical potential required to maintain a fixed $\langle N\rangle$ increases rapidly with $B$, in agreement with the spectrum $\varepsilon_{n}^{(3)}\propto B^{3/2}\sqrt{n(n-1)(n-2)}$ of ABC-stacked trilayer graphene. This sharp increase arises from the high density of low-energy states and the prominent degeneracy of the zero-energy levels.
Overall, these results clarify the dependence of the particle number and chemical potential on magnetic-field strength, providing the basis for the consistent implementation of thermodynamic cycles such as the Stirling cycle, examined in the next sections.
The analysis of the energy scales and the behavior of the chemical potential in monolayer, bilayer, and trilayer graphene reveals marked differences among these systems. To enable a consistent comparison, in this work the chemical potential is adjusted as a function of temperature and magnetic field such that the particle number remains fixed:
$N_{e}=\expectationvalue{N(\mu(B,T),B,T)}=\text{constant}.$ (19)
In this manner, the particle number $N_{e}$ in Eq. (19) is conserved throughout the entire thermodynamic cycle, with $N_{e}$ to determine $\mu(T,B)$ within the grand–canonical framework. This choice enables a direct comparison of the distinct scaling behaviors in mono-, bi-, and trilayer graphene, thereby highlighting the physical differences that originate from their respective Landau-level structures. In this framework, the chemical potential is effectively parametrized by enforcing particle-number conservation. Applying the Maxwell relation
$\left.\frac{d\mu}{dT}\right|_{N_{e}}=-\left.\frac{dS}{dN_{e}}\right|_{T}=0.$ (20)
It follows that the parametrized $\mu$ remains constant under temperature variations, while being nontrivially modified by the external magnetic field, as illustrated in Fig. 2.
For reference to experimental scales, we provide an order–of–magnitude estimate obtained from the magnetic length, $\ell_{B}=\sqrt{\hbar/(eB)}$. If $N_{e}$ electrons are assumed to occupy an effective area $\pi\ell_{B}^{2}$, the corresponding sheet density can be approximated as
@@
For instance, at $B=1~\mathrm{T}$ the estimated densities are $n_{2\mathrm{D}}\approx\{2.43,\,3.89,\,7.29\}\times 10^{11}~\mathrm{cm^{-2}}$ for $N_{e}=\{5,8,15\}$. This correspondence is not meant to represent an exact relation but rather to serve as a practical benchmark linking the effective particle number $N_{e}$ employed in the calculations with typical carrier densities reported in graphene samples.
Figure 3: Internal energy $U(T,B)$ as a function of temperature for (a) monolayer, (b) bilayer, and (c) trilayer graphene at different magnetic fields, with the particle number fixed at $N_{e}=15$.
Having established the Landau-level spectra in Eqs. 1, 3, and 5, together with the parametrized chemical potential of Eq. 19 that ensures a fixed particle number across all three systems. We now turn to the analysis of thermodynamic quantities directly relevant for the operation of quantum heat engines. In particular, we investigate how each multilayer configuration responds under thermal cycling, an aspect of central importance for the implementation of the Stirling cycle. Fig. 3 presents the internal energy of the monolayer, bilayer, and trilayer graphene as a function of temperature, for several fixed magnetic-field values and at constant particle number $N_{e}=15$.
@@ III.3 Thermodynamic Properties: Internal Energy and Entropy of Multilayer Graphene
Having established the Landau-level spectra in Eqs.(1, 3, 5) and the parametrized chemical potential in Eq.(19), we now turn to the analysis of thermodynamic quantities directly relevant for the operation of quantum heat engines. In particular, we examine how each multilayer configuration responds under thermal cycling, which is especially pertinent for the implementation of the Stirling cycle. The internal energy $U(T,B)$ of monolayer, bilayer, and trilayer graphene is presented in Fig. 3 as a function of temperature for several fixed magnetic-field values, at constant particle number $N_{e}=15$.
Figure 4: Entropy $S(T,B)$ as a function of temperature for (a) monolayer, (b) bilayer, and (c) trilayer graphene at different magnetic fields, with the particle number fixed at $N_{e}=15$.
@@
The next discussion focuses on the entropy to further characterize the thermodynamic response. Fig. 4 presents the temperature dependence of the entropy $S(T,B)$ for monolayer, bilayer, and trilayer graphene at different magnetic fields, with the particle number fixed at $N_{e}=15$. While all three systems exhibit a monotonic increase of entropy with temperature, notable differences arise in magnitude, curvature, and sensitivity to the magnetic field, underscoring the distinct thermodynamic responses of each multilayer configuration. In Fig. 4(a), the entropy of monolayer graphene remains negligible at low temperatures and begins to rise appreciably only above $T\simeq 25\,\mathrm{K}$, where thermal excitations populate higher Landau levels. The increase is markedly nonlinear, particularly at lower magnetic fields. This behavior reflects the relativistic Landau-level structure where stronger magnetic fields enlarge the spacing between energy levels and thereby suppress thermal activation. Consequently, higher fields substantially reduce the entropy and produce a clear separation among the curves.
@@ III.3.1 Quantum Stirling Cycle
Figure 5: Diagram of the quantum Stirling cycle in terms of entropy $S$ and the perpendicular magnetic field $B$. Curves AB and CD correspond to the isotherms at $T=T_{H}$ and $T=T_{L}$, respectively. As the magnetic field increases at fixed temperature, the entropy decreases, and vice versa.
The quantum Stirling cycle, illustrated in Fig. 5, consists of four strokes: two isothermal and two isomagnetic processes. In this framework, the perpendicular magnetic field $B$ acts as the external control parameter, while the working substance is assumed to remain in thermal equilibrium with the reservoirs during all stages of the cycle. The sequence of strokes is as follows:
@@
Table 1 summarizes the adopted sign convention for heat and total work across the four operational regimes considered. In this convention, a positive total work $W>0$ indicates that the system delivers mechanical work to the surroundings, whereas $W<0$ corresponds to work being supplied to the system. Similarly, a positive heat $Q>0$ denotes energy flowing into the working substance, while $Q<0$ represents heat released to the external reservoirs. Under these definitions, an engine operates with $W>0$, absorbing heat from the hot reservoir $(Q_{\mathrm{in}}>0)$ and rejecting heat to the cold one $(Q_{\mathrm{out}}<0)$.
@@
Beyond these two classical regimes, previous studies have shown that the quantum Stirling cycle may also operate in two additional regimes that arise purely from its quantum character Rastegar-Sedehi and Cruz (2024); Castorene et al. (2025b); Rastegar-Sedehi and Cruz (2025); Castorene et al. (2024); Araya et al. (2023); Behrendt and Deffner (2025). In the accelerator regime, both the total work and the heat released to the cold reservoir are negative, indicating that external work input is accompanied by simultaneous heating of both reservoirs. In the heater regime, all heat flows are negative and $W<0$, so that the external work is entirely dissipated as heat into both reservoirs. As will be shown in later sections, the exotic thermodynamic properties of the system, which arise from the constraint of particle number conservation, can induce these additional operational regimes.
@@ IV.1 Stirling Engine Efficiency
In the monolayer case shown in Fig. 6(a), the conditions for the system to operate as an engine are strongly constrained for thermal reservoirs at $T_{L}=50.7\,\mathrm{K}$ and $T_{H}=53.85\,\mathrm{K}$, corresponding to a Carnot efficiency of $\eta_{C}\simeq 0.06$, with the high magnetic field fixed at $B_{H}=4.0\,\mathrm{T}$. Although some curves (e.g., $N_{e}=5$) can, in principle, achieve the Carnot limit with finite work, this only occurs at low carrier densities and within narrow $B_{L}$ intervals. Outside these regions, the device switches to other operational regime modes and the definition of efficiency is not well defined. It is for this reason that only small intervals where the substance operates as an engine exist, while no efficiency curves are observed beyond them.
This restricted behavior can be interpreted as a consequence of the large spacing between relativistic Landau levels, which limits the thermal accessibility of higher-energy states. As a result, the heat exchanged during the isothermal strokes, while retaining the correct sign for engine operation, becomes smaller in magnitude. When the work contribution from the isomagnetic strokes surpasses that of the isothermal ones—an effect enhanced by the Landau level separation—the net work output turns negative and the cycle ceases to function as an engine. These exotic behaviors associated with monolayer graphene will be discussed in later sections.
For the bilayer graphene in Fig. 6(b), the system operates as an engine across the entire $B_{L}$ range considered, with the high magnetic field fixed at $B_{H}=3.18\,\mathrm{T}$. The thermal reservoirs are maintained at $T_{L}=45.9\,\mathrm{K}$ and $T_{H}=47.7\,\mathrm{K}$, corresponding to a Carnot efficiency of $\eta_{C}\simeq 0.04$. Under these conditions, the curves for $N_{e}=5$ and $N_{e}=15$ attain the Carnot bound at finite values of $B_{L}$, with the $N_{e}=15$ case attaining $\eta_{C}$ at lower $B_{L}$ values than $N_{e}=5$. In contrast, the $N_{e}=8$ curve remains slightly below $\eta_{C}$ over the full range, exhibiting a broad maximum just under the Carnot line at $B_{L}\simeq 0.8$[T]. This reflects a more balanced competition between the isothermal and isomagnetic contributions arising from the parametrization of the chemical potential. This intermediate behavior highlights how the bilayer spectrum, with its linear scaling in $B$, provides a broader operational window: depending on the carrier number, the cycle can either saturate the Carnot limit or remain close to it with finite work output.
@@ IV.3 Stirling Operational Regimes
In the previous sections, the working substances were shown to operate as engines with high efficiency and significant work extraction. Nevertheless, these systems may also display exotic behaviors, leading to the emergence of different operational regimes. To illustrate this point, Fig. 9 presents the operational map of the Stirling cycle for monolayer graphene at fixed $B_{H}=3.0~\mathrm{T}$, $T_{L}=46~\mathrm{K}$, and $N_{e}=5$. Within the explored window of low magnetic field $B_{L}$ and hot reservoir temperature $T_{H}$, no engine region is observed. Instead, the diagram reveals a small refrigerator domain at $B_{L}\simeq 0.7{-}1.8~\mathrm{T}$ and $T_{H}\simeq 46{-}49.5~\mathrm{K}$, two accelerator wedges near the lowest and highest values of $B_{L}$, and a broad heater region covering most of the parameter space.
@@
The absence of engine operation indicates that, under these conditions, heat exchange in the isomagnetic strokes dominates over the isothermal contributions. This outcome reflects both the wide spacing of relativistic Landau levels and the parametrization of the chemical potential imposed to maintain $N_{e}=5$ throughout the cycle. Consequently, the system exhibits $W<0$ across the entire diagram, with the specific regime determined by the balance of heat flows: refrigerator, accelerator, or heater. These results align with the performance maps in Fig. 8, where the monolayer shows weak $\eta W$ values and a rapidly vanishing optimum, as well as with the efficiency curves in Fig. 6, which display only small, low-efficiency engine regions and predominantly non-engine behavior. Engine windows may emerge under different parameter choices, but are absent in the present configuration.
@@
It should be emphasized that realizing different operational regimes in bilayer and trilayer graphene would require substantially lower temperatures and much higher magnetic fields than those explored in this study, owing to their smaller characteristic energy scales. Accordingly, the present analysis focuses on regimes accessible at comparatively higher temperatures, which are not only more relevant for the purposes of this work but also more feasible from an experimental perspective.

### v2

@@ III.2 Energy Scales and Parametrization of the Chemical Potential
For reference to experimental scales, we provide an estimate obtained from the magnetic length, $\ell_{B}=\sqrt{\hbar/(eB)}$ Castro Neto et al. (2009); Goerbig (2011). By defining $N_{e}$ as the number of electrons within the magnetic length area $\pi\ell_{B}^{2}$, the sheet density can be written as
@@
[…]
Fig. 2 provides an overview of the relation between the magnetic field $B$, the chemical potential $\mu$, and the resulting particle number $\langle N\rangle$ for mono-, bi-, and trilayer graphene at fixed temperature of $T=50$ K . In this figure, $N$ is not fixed; instead, we map the full surface $\langle N(50,B,\mu)\rangle$ in order to illustrate how the distinct Landau–level structures of each multilayer system shape the dependence of the particle number on $\mu$ and $B$. This exploration helps visualize the scaling behaviors originating from the linear, quadratic, and cubic dispersions of the three stackings.
To make a consistent comparison across all layers, the chemical potential is determined self-consistently from the particle-number constraint
$N_{e}=\frac{\left\langle N(T,B,\mu(T,B))\right\rangle}{\pi\ell_{B}^{2}}=\text{constant},$ (18)
so that $\mu(T,B)$ is obtained at every point of the thermodynamic cycle. Within this framework, the fixed particle number $N_{e}$ fully determines the temperature and magnetic-field dependence of the chemical potential, and no additional analytical assumption regarding its $T$-dependence is imposed.
We emphasize that, for fixed $B$, the chemical potential also acquires a (comparatively weaker) dependence on temperature, i.e., $\mu=\mu(T,B)$ is determined self-consistently throughout the cycle. This temperature dependence is fully included in our numerical evaluation of the thermodynamic quantities.
@@ III.3 Thermodynamic Properties: Internal Energy and Entropy of Multilayer Graphene
Figure 3: Internal energy $U(T,B,\mu(T,B))$} as a function of temperature for (a) monolayer, (b) bilayer, and (c) trilayer graphene at different magnetic fields, with the particle number fixed at $N_{e}=15$.
Having established the Landau-level spectra in Eqs. 4, 5, and 6, together with the parametrized chemical potential of Eq. 18 that ensures a fixed particle number across all three systems. We now turn to the analysis of thermodynamic quantities directly relevant for the operation of quantum heat engines. In particular, we investigate how each multilayer configuration responds under thermal cycling, an aspect of central importance for the implementation of the Stirling cycle.
The internal energy $U(T,B,\mu(T,B))$} of monolayer, bilayer, and trilayer graphene is presented in Fig. 3 as a function of temperature for several fixed magnetic-field values, at constant particle number $N_{e}=15$.
Figure 4: Entropy $S(T,B,\mu(T,B))$} as a function of temperature for (a) monolayer, (b) bilayer, and (c) trilayer graphene at different magnetic fields, with the particle number fixed at $N_{e}=15$.
@@
The next discussion focuses on the entropy to further characterize the thermodynamic response. Figure 4 presents the temperature dependence of the entropy $S(T,B,\mu(T,B))$} for monolayer, bilayer, and trilayer graphene at different magnetic fields, with the particle number fixed at $N_{e}=15$. While all three systems exhibit a monotonic increase of entropy with temperature, notable differences arise in magnitude, curvature, and sensitivity to the magnetic field, underscoring the distinct thermodynamic responses of each multilayer configuration. In Fig. 4(a), the entropy of monolayer graphene remains negligible at low temperatures and begins to rise appreciably only above $T\simeq 25\,\mathrm{K}$, where thermal excitations populate higher Landau levels. The increase is markedly nonlinear, particularly at lower magnetic fields. This behavior reflects the relativistic Landau-level structure where stronger magnetic fields enlarge the spacing between energy levels and thereby suppress thermal activation. Consequently, higher fields substantially reduce the entropy and produce a clear separation among the curves.
@@ III.3.1 Quantum Stirling Cycle
Figure 5: Diagram of the quantum Stirling cycle in terms of entropy $S(T,B,\mu(T,B))$ for fixed particle-number $N_{e}$ and the perpendicular magnetic field $B$. Curves AB and CD correspond to the isotherms at $T=T_{H}$ and $T=T_{L}$, respectively. As the magnetic field increases at fixed temperature, the entropy decreases, and vice versa.
The quantum Stirling cycle, illustrated in Fig. 5, consists of four strokes: two isothermal and two isomagnetic processes. In this framework, the perpendicular magnetic field $B$ acts as the external control parameter, while the working substance remains in thermal equilibrium with the reservoirs throughout the entire cycle. For each pair of values $(T,B)$, the chemical potential $\mu(T,B)$ is determined self-consistently to keep the particle number $N_{e}$ fixed, and the resulting value is then used in all thermodynamic quantities. The sequence of strokes is as follows:
@@
Table 1 summarizes the adopted sign convention for heat and total work across the four operational regimes considered. In this convention, a positive total work $W>0$ indicates that the system delivers work to the surroundings, whereas $W<0$ corresponds to work being supplied to the system. Similarly, a positive heat $Q>0$ denotes energy flowing into the working substance, while $Q<0$ represents heat released to the external reservoirs. Under these definitions, an engine operates with $W>0$, absorbing heat from the hot reservoir $(Q_{\mathrm{in}}>0)$ and rejecting heat to the cold one $(Q_{\mathrm{out}}<0)$.
@@
Beyond these two classical regimes, previous studies have shown that the quantum Stirling cycle may also operate in two additional regimes that arise purely from its quantum character Rastegar-Sedehi and Cruz (2024); Castorene et al. (2025b); Rastegar-Sedehi and Cruz (2025); Castorene et al. (2024); Araya et al. (2023); Behrendt and Deffner (2025). In the accelerator regime, both the total work and the heat released to the cold reservoir are negative, indicating that external work input is accompanied by simultaneous heating of both reservoirs. In the heater regime, all heat flows are negative and $W<0$, so that the external work is entirely dissipated as heat into both reservoirs.
Importantly, these operational regimes are not exclusive to intrinsic properties of stacked graphene systems. It has been shown that twisted bilayer graphene, when employed as a Stirling engine at temperatures within 50 to 150 K, can also exhibit this diversity of thermodynamic behaviors Soufy and Benjamin (2025). As will be shown in later sections, the exotic thermodynamic properties of the system, which arise from the constraint of particle number conservation, can induce these additional operational regimes.
@@ IV.1 Stirling Engine Efficiency
In the monolayer case shown in Fig. 6(a), the conditions for the system to operate as an engine are strongly constrained for thermal reservoirs at $T_{L}=50.7\,\mathrm{K}$ and $T_{H}=53.85\,\mathrm{K}$, corresponding to a Carnot efficiency of $\eta_{C}\simeq 0.06$, with the high magnetic field fixed at $B_{H}=4.0\,\mathrm{T}$. Although some curves (e.g., $N_{e}=5$) can, in principle, achieve the Carnot limit with finite work, this only occurs at low carrier densities and within narrow $B_{L}$ intervals. Outside these regions, the device switches to other operational regime modes, and the definition of efficiency is not well-defined. It is for this reason that only small intervals where the substance operates as an engine exist, while no efficiency curves are observed beyond them.
This restricted behavior can be interpreted as a consequence of the large spacing between relativistic Landau levels, which limits the thermal accessibility of higher-energy states. As a result, the heat exchanged during the isothermal strokes, while retaining the correct sign for engine operation, becomes smaller in magnitude. When the heat contribution from the isomagnetic strokes surpasses that of the isothermal ones—an effect enhanced by the Landau level separation—the net work output turns negative, and the cycle ceases to function as an engine. These exotic behaviors associated with monolayer graphene will be discussed in later sections.
For the bilayer graphene in Fig. 6(b), the system operates as an engine across the entire $B_{L}$ range considered, with the high magnetic field fixed at $B_{H}=3.18\,\mathrm{T}$. The thermal reservoirs are maintained at $T_{L}=45.9\,\mathrm{K}$ and $T_{H}=47.7\,\mathrm{K}$, corresponding to a Carnot efficiency of $\eta_{C}\simeq 0.04$. Under these conditions, the curves for $N_{e}=5$ and $N_{e}=15$ attain the Carnot bound at finite values of $B_{L}$, with the $N_{e}=15$ case attaining $\eta_{C}$ at lower $B_{L}$ values than $N_{e}=5$. In contrast, the $N_{e}=8$ curve remains slightly below $\eta_{C}$ over the full range, exhibiting a broad maximum just under the Carnot line at $B_{L}\simeq 0.8$[T]. This reflects a more balanced competition between the isothermal and isomagnetic contributions arising from the parametrization of the chemical potential. This intermediate behavior highlights how the bilayer spectrum, with its linear scaling in $B$, provides a broader operational window: depending on the carrier number, the cycle can either saturate the Carnot limit or remain close to it with finite work output. This behavior contrasts with results for twisted bilayer graphene studied within the canonical ensemble Soufy and Benjamin (2025). While such systems can reach Carnot efficiency in Otto cycles Singh and Benjamin (2021) and exhibit high efficiencies in quantum Stirling cycles, saturation of the Carnot bound in the Stirling cycle has not been reported in that framework. This suggests that a Fermi–Dirac description with fixed particle number enhances thermodynamic performance in graphene-based quantum heat engines, highlighting the role of ensemble choice and carrier-number constraints in setting efficiency limits.
@@ IV.3 Stirling Operational Regimes
In the previous sections, the working substances were shown to operate as engines with high efficiency and significant work extraction. Nevertheless, these systems may also exhibit exotic behavior, leading to the emergence of distinct operational regimes. To illustrate this point, Fig. 9 presents the operational map of the Stirling cycle for monolayer graphene at fixed $B_{H}=3.0~\mathrm{T}$, $T_{L}=46~\mathrm{K}$, and $N_{e}=5$. Within the explored window of low magnetic field $B_{L}$ and hot reservoir temperature $T_{H}$, no engine region is observed. Instead, the diagram reveals a small refrigerator domain at $B_{L}\simeq 0.7{-}1.8~\mathrm{T}$ and $T_{H}\simeq 46{-}49.5~\mathrm{K}$, two accelerator wedges near the lowest and highest values of $B_{L}$, and a broad heater region covering most of the parameter space.
@@
The absence of engine operation indicates that, under these conditions, heat exchange in the isomagnetic strokes dominates over the isothermal contributions. This outcome reflects both the wide spacing of relativistic Landau levels and the parametrized chemical potential imposed to maintain $N_{e}=5$ throughout the cycle. Consequently, the system exhibits $W<0$ across the entire diagram, with the specific regime determined by the balance of heat flows: refrigerator, accelerator, or heater. These results align with the performance maps in Fig. 8, where the monolayer shows weak $\eta W$ values and a rapidly vanishing optimum, as well as with the efficiency curves in Fig. 6, which display only small, low-efficiency engine regions and predominantly non-engine behavior. Engine windows may emerge under different parameter choices, but are absent in the present configuration.
This does not exclude bilayer graphene from exhibiting the full range of operational regimes. Previous studies have shown that the same Stirling cycle implemented in bilayer graphene can display all operational regimes under different parameter choices and when the chemical potential is not fixed Soufy and Benjamin (2025). Their absence here therefore reflects the specific thermodynamic window considered, rather than an intrinsic limitation of the system.
@@
It should be emphasized that realizing different operational regimes in bilayer and trilayer graphene would require substantially lower temperatures and much higher magnetic fields than those explored in this study, owing to their smaller characteristic energy scales. Accordingly, the present analysis focuses on regimes accessible at higher temperatures, which are not only more relevant to this work but also more feasible from an experimental perspective.
At the same time, it is important to note that the present analysis is intended as a theoretical proof of concept rather than a concrete experimental proposal. While the main text focuses on a fixed-particle-number formulation in order to enable a controlled comparison between different multilayer systems, an alternative and experimentally natural realization of the Stirling cycle consists in fixing the chemical potential. This extension, which mimics the situation of a graphene system coupled to electronic reservoirs, is explicitly addressed in Appendix A, where we show that the qualitative features of the operational regimes remain robust with respect to the choice of thermodynamic ensemble.
In a realistic implementation, field-driven modifications of the Landau–level structure would induce variations of the sample magnetization across the different strokes of the cycle. When the graphene system is electromagnetically coupled to an external circuit, these magnetization changes can generate an electromotive response, allowing the conversion of field-induced energy variations into usable electrical work. This mechanism provides a physically consistent route for work extraction without requiring commitment to a specific device architecture.